# New Foundations is consistent

### v22

@@ A $-1$-approximation is a function $\psi$ such that:
Lemma 4.6 (adding orbits).
Let $\psi$ be a $-1$-approximation, and let $(L_{n})_{n\in\mathbb{Z}}$ be litters such that
$\{\langle L_{n},L_{n+1}\rangle:n\in\mathbb{Z}\}$
is a bijection, and $\psi^{*}$ is not defined at any of the $L_{n}$. Then $\psi$ has an extension $\chi$ where $\chi^{*}``L_{n}\sim L_{n+1}$ for each $n$, and all near-litters contained in the domain of $\chi^{*}$ but not contained in the domain of $\psi^{*}$ are near some $L_{n}$.
Proof.
Define
$\chi=\psi\cup\{\langle L_{n},L_{n+1}\rangle:n\in\mathbb{Z}\};$
it is easy to verify all of the required conditions. ∎
It is not necessary that all of the $L_{n}$ be distinct. This lemma therefore allows us to add single orbits of litters of any order to $-1$-approximations.
@@
    If $A$ is a finite subset of $\lambda\cup\{-1\}$ with maximum element $\beta$, we define the derivative of $\psi$ along $A$ in the following way: $\psi_{A}=(\psi_{A\setminus\{{\tt min}(A)\}})_{{\tt min}(A)}$ and $\psi_{\{\beta\}}=\psi$. If $\beta=-1$, construe $\psi_{-1}$ as $\psi$.
@@ A $\beta$-approximation $\psi$ acts on a support $S$ by
    If $L$ is $A$-flexible for some $A$, then $\psi_{A}^{*}``L$ is also $A$-flexible.
@@
    $\psi_{A}^{*}``f_{\delta,\gamma}(x,S)\sim f_{\delta,\gamma}(\pi(x),\pi[S])$
    (and hence all $\delta$-allowable permutations $\pi$ satisfying the given hypotheses also satisfy the stated coherence condition).
@@
If $\psi$ is coherent, then $\psi^{n}$ is coherent for any integer $n$, and $\psi_{\gamma}$ is coherent for any $\gamma<\beta$. Every $-1$-approximation is coherent.
@@
Let $\psi$ be a coherent $\beta$-approximation. Let $A$ be an extended type index with maximum $\beta$, and let $\gamma={\tt min}(A_{1})$ and $\delta<{\tt min}(A_{2})$. Let $(x,S)\in\tau_{\delta}^{+}$ be such that $(\psi_{A_{2}})_{\delta}^{*}[S]$ is defined. Then if freedom of action holds at type level $\delta$, there is a coherent extension $\chi\succeq_{\beta}\psi$ such that $\chi_{A}^{*}``f_{\delta,\gamma}(x,S)$ is defined.
@@
Otherwise, we can use lemma 4.6 to extend $\psi$ to an approximation $\chi$ in which
@@
for each integer $n$. It is easy to check that this $\chi$ is coherent. ∎
@@
Suppose that $\chi_{A}^{*}$ is not defined on all litters for some $A$. Let $L$ be the litter with minimal position $\iota_{*}(L)$ such that there is an extended type index $A$ with maximum element $\beta$ such that $L\nsubseteq{\tt dom}(\psi_{A}^{*})$.
@@
Now suppose that $L$ is not $A$-flexible. Writing $\gamma$ for the minimum element of $A$, there are $\delta<{\tt min}(A_{1})$ and $(x,S)\in\tau_{\delta}^{*}$ such that $L=f_{\delta,\gamma}(x,S)$.
We claim that $(\psi_{A_{2}})_{\delta}^{*}[S]$ is defined; this will then give a contradiction by lemma 4.17. To show this, we must prove that if $(t,B)\in{\tt rng}(S)$, then $\psi_{A}^{*}``t$ is defined. By definition 3, we have $\iota_{*}^{+}(x,S)<\iota_{*}(L)$, and by (I11), $\iota_{*}(t)\leq\iota_{*}^{+}(x,S)$. If $t$ is a near litter or singleton of an atom and $t\subset M$ for a litter $M$, then by definition 3.10, we obtain $\iota_{*}(M)<\iota_{*}(t)$, and therefore that $\psi_{B}^{*}``t$ is defined by minimality of the position of $L$.Alternatively, if $t$ is a near-litter, then if $M$ is any litter such that $M\cap t\neq\emptyset$, definition 3.10 again implies that $\iota_{*}(M)<\iota_{*}(t)$, so $\psi_{B}^{*}``M$ is defined. Combining all such litters, we conclude that $\psi_{B}^{*}``t$ is defined.
Therefore $\chi_{A}^{*}$ must be defined on all litters for all $A$. By remark 4.14, this concludes the proof: $\psi$ approximates the allowable permutation $\pi$ given by $\pi_{A}^{*}=\chi_{A}^{*}$. ∎
@@
In fact, all of the proofs of this subsection can be phrased in such a way that the axiom of choice is not invoked. In particular, the coherent extension defined in 4.17 is uniquely determined: all choices of $\pi$ yield the same extension $\chi$. This means that every coherent approximation defined on all flexible litters has a unique coherent extension defined on all litters (but we will not use this fact).
4.2 Types are of size $\mu$ (so the construction actually succeeds)
@@ A $\beta$-support $S$ is called strong if it satisfies the additional properties that
    if $(x,A)$ and $(y,A)$ are in the range of $S$, then ($(\{z\},A)\in S$ for all atoms $z$ in the interference of $x$ and $y$,
@@
    and for each $\epsilon$ in the domain of $S$, if $S_{\epsilon}=(x,A)$, and $x^{\circ}=f_{\gamma,\delta}(y,T)$, then the range of $T^{\uparrow A_{2}}$ is a subset of the range of $S\lceil\epsilon$: supports appearing in inverse images under $f$ of litters which are near the first projections of an element of the support have a type-raised copy (mod reindexing) appearing in the support before that item.
@@
Each support $S$ is the terminal segment of a strong support. Such a strong support can be constructed by prefixing to $S$, for each $T^{\uparrow A_{2}}$ such that for some $\epsilon$ in the domain of $S$, $S_{\epsilon}=(x,A)$, and $x^{\circ}=f_{\gamma,\delta}(y,T)$, a strong support with $T^{\uparrow A_{2}}$ as a terminal segment, which will be obtainable as $U^{\uparrow A_{2}}$ where $U$ is a strong support with $T$ as a terminal segment, which exists by the inductive hypothesis that this is true for supports with lower type index than $S$, followed by all atomic items which must be added to satisfy the first condition in the definition of strong support. In fact, we can define a canonical strong support of which $S$ is a terminal segment by stipulating that the $U^{\uparrow A_{2}}$’s are added (each one in turn between the preceding ones and $S$) in the order in which the correlated $S_{\epsilon}$’s appear in $S$ and that $U$ is the canonical downward extension of $T$ in each case, and that the atomic items are added in the order of their images under $\iota^{+}_{*}$, with some fixed well-ordering of extended type indices used to resolve order of items with the same value under $\iota^{+}_{*}$.
Definition 4.24 (coding function).
For any support $S$ and object $x$, we can define a function $\chi_{x,S}$ which sends $T=\pi[S]$ to $\pi(x)$ for every $T$ in the orbit of $S$ under the action of allowable permutations. We call such functions coding functions. Note that if $\pi[S]=\pi^{\prime}[S]$ then $(\pi^{-1}\circ\pi^{\prime})[S]=S$, so $(\pi^{-1}\circ\pi^{\prime})(x)=x$, so $\pi(x)=\pi^{\prime}(x)$, ensuring that the map $\chi_{x,S}$ for which we gave an implicit definition is well defined.
Definition 4.25 (designated support).
For each ordinal $\gamma$, and for each orbit in $\tau_{\gamma}$ under allowable permutations, choose $x$ in the orbit (the designated element of the orbit), choose a strong support $S$ of $x$, and for every other $y$ in the orbit, choose a designated allowable permutation $\pi_{y}$ such that $\pi_{y}(x)=y$, and define the designated support of $y$ to be $\pi_{y}[S]$.
Definition 4.26 (specification).
A $\eta$-specification $S^{*}$ of a $\eta$-support $S$ is a function with the same domain as $S$. We use the notation $S^{*}_{\epsilon}$ for $S^{*}(\epsilon)$.
  1. 1.
    If $S_{\epsilon}$ is ($(\{x\},A)$, where $\beta={\tt min}(A_{1})$, then $S^{*}_{\epsilon}$ is $(0,\beta,\Sigma,A)$ where $\Sigma$ is the set of all $\delta$ such that $A=\pi_{2}(S_{\delta})$ and $x\in\pi_{1}(S_{\delta})$ (this captures identical atoms and near litters containing the given atom)
  2. 2.
    If $S_{\epsilon}$ is $(N,A)$ and $N$ is a near-litter, where $\beta={\tt min}(A_{1})$, and either $|A|\leq 2$ or $N^{\circ}$ is not in the range of any $f_{\gamma,\beta}$ for $\gamma<{\tt min}(A_{2})$ (that is, $N$ is $A$-flexible), then $S^{*}_{\epsilon}$ is $(1,\beta,\Sigma,A)$, where $\Sigma$ is the set of all $\delta$ such that $\pi_{2}(S_{\delta})=A$ and $\pi_{1}(S_{\delta})\sim\pi_{1}(S_{\epsilon})$.
  3. 3.
    If $S_{\epsilon}$ is $(N,A)$ and $N$ is a near-litter, where $\beta={\tt min}(A_{1})$, and $N^{\circ}=f_{\gamma,\beta}(x,T)$ with $-1\leq\gamma<{\tt min}(A_{2})$ and $x\in\tau_{\gamma}$ then $S^{*}_{\epsilon}$ is $(2,\beta,\chi_{x,T},F,A)$, where $F$ is a function from the domain of $T$ into $\epsilon$ such that $S_{F(\delta)}=(T^{\uparrow A_{2}})_{\delta}$ for each $\delta$ in the domain of $T$, or 1 if there is no such $F$ (the usefulness of 1 as a dummy being that it is not a function). There is a method to choose a canonical such $F$ if there is one: add the provision that for each $\delta$, $F(\delta)$ is chosen as small as possible.
Remark 4.27.
@@
It should be evident that every support has a specification, and that a strong support will have a specification with no instances of $F=1$, and that for any $\beta$-allowable permutation $\pi$ and strong $\beta$-support $S$, $(\pi[S])^{*}=S^{*}$. What is less evident and our first target result here is that if $S$ is a strong support then any $T$ with $T^{*}=S^{*}$ is the image of $S$ under the action of an allowable permutation: the specifications precisely code the orbits in the strong supports under the allowable permutations.

### v23

@@ 4.1 The Freedom of Action theorem
In this section, we will prove a theorem that gives us a way to construct allowable permutations (theorem 4.17). To do this, we will define a certain class of objects for each type, called approximations, that behave like partial structural permutations. We will say that an approximation is called coherent if it satisfies a particular version of the coherence condition on allowable permutations (definition 3.20). We will then show that every coherent approximation can be identified with the restriction of some allowable permutation.
@@ A $-1$-approximation is a function $\psi$ such that:
    If $A$ is a finite subset of $\lambda\cup\{-1\}$ with maximum element $\beta$, we define the derivative of $\psi$ along $A$ in the following way: $\psi_{A}=(\psi_{A\setminus\{{\tt min}(A)\}})_{{\tt min}(A)}$ and $\psi_{\{\beta\}}=\psi$.
    If $\beta=-1$, construe $\psi_{-1}$ as $\psi$.
@@ A $\beta$-approximation $\psi$ acts on a support $S$ by
    If $L$ is $A$-flexible for some $A$, then $\psi_{A}^{*}``L$ is also $A$-flexible whenever this is defined.
@@
    $\psi_{A}^{*}``f_{\delta,\gamma}(x,S)\sim f_{\delta,\gamma}(\pi(x),\pi[S]).$
    Note that in this case, if $\pi^{\prime}$ is any other $\delta$-allowable permutation satisfying
    $(\psi_{A_{2}})_{\delta}^{*}[S]=\pi^{\prime}[S]$
    then we can conclude
    $\psi_{A}^{*}``f_{\delta,\gamma}(x,S)\sim f_{\delta,\gamma}(\pi^{\prime}(x),\pi^{\prime}[S])$
    as $\pi$ and $\pi^{\prime}$ have actions that coincide on $S$ and therefore at $x$.
@@
If $\psi$ is coherent, then $\psi^{n}$ is coherent for any integer $n$, and $\psi_{\epsilon}$ is coherent for any $\epsilon<\beta$. Every $-1$-approximation is coherent.
@@
Let $\psi$ be a coherent $\beta$-approximation. Let $A$ be an extended type index with maximum $\beta$, and let $\gamma={\tt min}(A_{1})$ and $\delta<{\tt min}(A_{2})$. Let $(x,S)\in\tau_{\delta}^{+}$ be such that $(\psi_{A_{2}})_{\delta}^{*}[S]$ is defined.
Then if freedom of action holds at type level $\delta$, there is a coherent extension $\chi\succeq_{\beta}\psi$ such that $\chi_{A}^{*}``f_{\delta,\gamma}(x,S)$ is defined.
@@
Otherwise, we can extend $\psi$ to an approximation $\chi$ given by
$\chi_{A}=\psi_{A}\cup\{(f_{\delta,\gamma}(\pi^{n}(x),\pi^{n}[S]),f_{\delta,\gamma}(\pi^{n+1}(x),\pi^{n+1}[S])):n\in\mathbb{Z}\}$
and $\chi_{B}=\psi_{B}$ for all $B\neq A$. It is easy to see that $\chi$ is an extension of $\psi$, and
@@
for each integer $n$. Finally, we must check that this $\chi$ is coherent, but this holds by construction. ∎
@@
Suppose that $\chi_{A}^{*}$ is not defined on all litters for some $A$. Let $L$ be the litter with minimal position $\iota_{*}(L)$ such that there is an extended type index $A$ with maximum element $\beta$ such that $L\nsubseteq{\tt dom}(\chi_{A}^{*})$.
@@
Now suppose that $L$ is not $A$-flexible. Writing $\gamma={\tt min}(A_{1})$, there are $\delta<{\tt min}(A_{2})$ and $(x,S)\in\tau_{\delta}^{+}$ such that $L=f_{\delta,\gamma}(x,S)$.
We claim that $(\chi_{A_{2}})_{\delta}^{*}[S]$ is defined; this will then give a contradiction by lemma 4.16. To show this, we must prove that if $(t,B)\in{\tt rng}(S)$, then $(\chi_{A_{2}})_{B}^{*}``t$ is defined. By definition 7, we have $\iota_{*}^{+}(x,S)<\iota_{*}(L)$, and by (I11), $\iota_{*}(t)\leq\iota_{*}^{+}(x,S)$. If $t$ is a singleton of an atom and $t\subset M$ for a litter $M$, then by definition 3.10, we obtain $\iota_{*}(M)<\iota_{*}(t)$, and therefore that $(\chi_{A_{2}})_{B}^{*}``t$ is defined by minimality of the position of $L$. Alternatively, if $t$ is a near-litter, then if $M$ is any litter such that $M\cap t\neq\emptyset$, definition 3.10 again implies that $\iota_{*}(M)\leq\iota_{*}(t)$, so $(\chi_{A_{2}})_{B}^{*}``M$ is defined. Combining all such litters, we conclude that $(\chi_{A_{2}})_{B}^{*}``t$ is defined.
Therefore $\chi_{A}^{*}$ must be defined on all litters for all $A$. By remark 4.13, this concludes the proof: $\psi$ approximates the allowable permutation $\pi$ given by $\pi_{A}=\chi_{A}^{*}$. ∎
@@
In fact, all of the proofs of this subsection can be phrased in such a way that the axiom of choice is not invoked. In particular, the coherent extension defined in 4.16 is uniquely determined: all choices of $\pi$ yield the same extension $\chi$. This means that every coherent approximation defined on all flexible litters has a unique coherent extension defined on all litters (but we will not use this fact).
4.2 Freedom of Action for incomplete orbits
In this subsection, we give an alternative form of the Freedom of Action Theorem (4.17) that is more useful when we do not already have full orbits of atoms or litters for use in approximations.
@@ 4.2 Freedom of Action for incomplete orbits
Definition 4.20 (actions).
A $-1$-action is a function $\xi$ such that:
  1. 1.
    The domain and image of $\xi$ are small, and $\xi$ is injective.
  2. 2.
    Each domain element of $\xi$ is either an atom or a near-litter, and moreover, $\xi$ maps atoms to atoms and near-litters to near-litters.
  3. 3.
    If $\xi$ is defined at near-litters $N_{1}$ and $N_{2}$, then $N_{1}^{\circ}=N_{2}^{\circ}$ if and only if $\xi(N_{1})^{\circ}=\xi(N_{2})^{\circ}$. This ensures that $\xi$ gives rise to an injective partial function on litters.
  4. 4.
    Whenever $a,N\in{\tt dom}(\xi)$, we have $a\in N$ if and only if $\xi(a)\in\xi(N)$.
  5. 5.
    The interference of domain elements of $\xi$ is included in the domain of $\xi$, and the interference of range elements of $\xi$ is included in the range of $\xi$.
Note that this definition ensures that the inverse of a $-1$-action is a $-1$-action.
Definition 4.21.
For $-1<\beta$, a $\beta$-action is a function with $\{-1\}\cup\beta$ as domain such that $\xi(\gamma)$ is a $\gamma$-action for each $\gamma<\beta$. For convenience, if $a$ is an atom we define $\xi_{A}(\{a\})$ as $\{\xi_{A}(a)\}$.
Definition 4.22.
We define derivatives for actions in the same way as we do for approximations. Actions can be applied to supports as follows:
$(\xi[S])_{\delta}=(\xi_{A}(x),A)\ \mathrm{where}\ S_{\delta}=(x,A)$
As with approximations, this operation is partial. An action is called coherent if:
  1. 1.
    If $N$ is $A$-flexible for some $A$, then $\xi_{A}(N)$ is also $A$-flexible whenever this is defined.
  2. 2.
    If $A$ is an extended type index with maximum $\beta$ and ${\tt min}(A_{1})=\gamma<\beta$, and $\delta<{\tt min}(A_{2})$ and $(x,S)\in\tau_{\delta}^{+}$ are such that $f_{\delta,\gamma}(x,S)\sim N\in{\tt dom}(\xi_{A})$, then there is some $\delta$-allowable permutation $\pi$ such that
    $(\xi_{A_{2}})_{\delta}[S]=\pi[S]\ \mathrm{and}\ \xi_{A}(N)\sim f_{\delta,\gamma}(\pi(x),\pi[S]).$
A $\beta$-action $\xi$ is said to approximate a $\beta$-allowable permutation $\pi$ if $\xi_{A}``x=\pi_{A}``x$ whenever $x$ is a singleton of an atom or a near-litter, and the left-hand side is defined.
Theorem 4.23 (Freedom of Action for incomplete orbits).
Every coherent $\beta$-action approximates some $\beta$-allowable permutation $\pi$.
Proof.
Let $\xi$ be a coherent $\beta$-action. First, note that we may assume without loss of generality that whenever $N\in{\tt dom}(\xi_{A})$, we have
$N\Delta N^{\circ}\subseteq{\tt dom}(\xi_{A})\ \mathrm{and}\ \xi_{A}(N)\Delta\xi_{A}(N)^{\circ}\subseteq{\tt rng}(\xi_{A}).$ (*)
We can extend any $\xi$ to an action $\xi^{\prime}$ with this property. If there is an atom in $N^{\circ}\setminus N\setminus{\tt dom}(\xi_{A})$ then neither $N^{\circ}$ nor indeed any $N^{\prime}\sim N$ containing the atom is in the domain of $\xi$, and we can map each such atom into an unused litter. If there is an atom in $N\setminus N^{\circ}\setminus{\tt dom}(\xi_{A})$ then again $N^{\circ}$ is not in the domain of $\xi$ and all $N^{\prime}\sim N$ which are in the domain of $\xi$ contain the atom, and we can map each such atom to an unused atom which is in $\xi_{A}(N)^{\circ}$ and also inside all $\xi_{A}$-images of near-litters $N^{\prime}\sim N$. This handles near-litters $N$ in the domain of $\xi$; near-litters in the range of $\xi$ are handled dually. This modification preserves coherence, as it only adds images of atoms.
Now, we will turn $\xi$ into a $\beta$-approximation $\psi$. To do this, we need to fill in orbits of litters and atoms.
We first show how to define orbits of flexible litters; the orbits for inflexible litters will appear automatically by coherence. For each $A$, let $g_{A}$ be the partial function given by $g_{A}(N^{\circ})=\xi_{A}(N)^{\circ}$ whenever $N$ is $A$-flexible and $\xi_{A}(N)$ is defined; the third condition on the definition of actions ensures that this map is well-defined and injective. We define the restriction of $\psi_{A}$ to litters to be a small bijective extension of $g_{A}$, where any new litters in ${\tt dom}(\psi_{A})$ are $A$-flexible and disjoint from domain and range elements of $\xi_{A}$.
Next, we show how to define orbits of atoms. For each $A$, we define the restriction of $\psi_{A}$ to atoms to be a small bijective extension of the restriction of $\xi_{A}$ to atoms, with the property that $a\in N$ iff $\psi_{A}(a)\in\xi_{A}(N)$ whenever the right-hand side is defined. We now explain how to do this. Our construction proceeds in $\omega$ stages, where at each stage, we define $\psi_{A}$ and $\psi_{A}^{-1}$ at any atoms contained in the domain or range of the function defined so far. We describe the construction of images under $\psi_{A}$; the inverse is dual.
Suppose that we need to define $\psi_{A}(a)$, so in particular $a\notin{\tt dom}(\xi_{A})$, and that $\xi_{A}(N)$ is defined for some $N\ni a$. By ()(*)( ∗ ), we know that $a\in N^{\circ}$, and additionally $a\in N^{\prime}$ for every $N^{\prime}\sim N$ with $N^{\prime}\in{\tt dom}(\xi_{A})$ because of the interference conditions. So we define $\psi_{A}(a)$ to be any unused atom both in $\xi_{A}(N)^{\circ}$ and also inside all $\xi_{A}(N^{\prime})$ for $N^{\prime}\sim N$. Such an atom will always exist because there are only a small amount of atoms excluded by each near-litter $N^{\prime}$, and there are only a small number of such near-litters.
If $\xi_{A}(N)$ is not defined for any $N\ni a$, then we are free to choose any unused atom for $\psi_{A}(a)$.
We perform the dual construction to define $\psi_{A}^{-1}$ on required atoms. When the construction concludes, the function $\psi_{A}$ will still be small because $\kappa$ is uncountable and regular.
The $\beta$-approximation $\psi$ defined as above is obviously coherent because all of the litters $L$ in the domain of $\psi_{A}$ are $A$-flexible, so by the Freedom of Action Theorem (4.17), $\psi$ approximates some $\beta$-allowable permutation $\pi$.
We claim that $\xi$ approximates this $\pi$. The case for atoms is obvious, so it suffices to show that for each near-litter $N$ such that $\xi_{A}(N)$ is defined, we have $\xi_{A}(N)=\pi_{A}``N$. By induction we may assume that $\xi_{B}(M)=\pi_{B}``M$ holds for all extended type indices $B$ and near-litters $M$ with $\iota_{*}(M)<\iota_{*}(N)$.
First, we show that $\xi_{A}(N)^{\circ}=(\pi_{A}``N)^{\circ}$. If $N$ is $A$-flexible, we have $\xi_{A}(N)^{\circ}=\psi_{A}(N^{\circ})=(\pi_{A}``N)^{\circ}$ as required. Suppose instead that $\gamma={\tt min}(A_{1})$ and $\delta<{\tt min}(A_{2})$ are such that $N\sim f_{\delta,\gamma}(x,S)$ for $(x,S)\in\tau_{\delta}^{+}$. For any $(t,B)$ in the range of $S$, by coherence of $\xi$ we know that $(\xi_{A_{2}})_{B}(t)$ is defined, and moreover by inductive hypothesis, that $(\xi_{A_{2}})_{B}(t)=(\pi_{A_{2}})_{B}``t$. But this shows that $\xi_{A_{2}}$ agrees with $\pi_{A_{2}}$ on $S$, and so again by coherence, we can conclude that
$\xi_{A}(N)\sim f_{\delta,\gamma}((\pi_{A_{2}})_{\delta}(x),(\pi_{A_{2}})_{\delta}[S])\sim\pi_{A}``f_{\delta,\gamma}(x,S)\sim\pi_{A}``N.$
We now conclude the proof by showing that $\xi_{A}(N)=\pi_{A}``N$. This proof depends on ()(*)( ∗ ), as well as the fact that the allowable permutation $\pi$ that we constructed in theorem 4.17 has the property that for any atom $a\notin{\tt dom}(\psi)$ and any litter $L$, we have $a\in L$ if and only if $\pi(a)\in(\pi``L)^{\circ}$.
Suppose that $a\in\xi_{A}(N)$. If $a\in{\tt rng}(\psi_{A})$, then $\psi_{A}^{-1}(a)\in N$ by construction of $\psi_{A}$, but $\psi_{A}^{-1}(a)=\pi_{A}^{-1}(a)$, so $\pi_{A}^{-1}(a)\in N$. Otherwise, $a\in\xi_{A}(N)^{\circ}$ by ()(*)( ∗ ), so $\pi_{A}^{-1}(a)\in N^{\circ}$, so $\pi_{A}^{-1}(a)\in N$ by ()(*)( ∗ ).
Conversely, suppose that $a\in\pi_{A}``N$. If $a\in{\tt rng}(\psi_{A})$, then $\pi_{A}^{-1}(a)=\psi_{A}^{-1}(a)$, so $a\in\xi_{A}(N)$ by construction of $\psi_{A}$. Otherwise, $\pi_{A}^{-1}(a)\in N^{\circ}$ by ()(*)( ∗ ), so $a\in(\pi_{A}``N)^{\circ}=\xi_{A}(N)^{\circ}$, so by ()(*)( ∗ ) we conclude that $a\in\xi_{A}(N)$. ∎
4.3 Counting orbits of supports: coding functions and specifications
@@ A $\beta$-support $S$ is called strong if it satisfies the additional properties that
    if $(x,A)$ and $(y,A)$ are in the range of $S$, then ($(\{z\},A)$ is in the range of $S$ for all atoms $z$ in the interference of $x$ and $y$ (defined in definition 4.19),
@@
    and if $(x,A)$ is in the range of $S$, $\delta={\tt min}(A_{1})$, $\gamma<{\tt min}(A_{2})$, and $x^{\circ}=f_{\gamma,\delta}(y,T)$, then the range of $T^{\uparrow A_{2}}$ is a subset of the range of $S$: supports appearing in inverse images under $f$ of litters which are near the first projections of an element of the support have a type-raised copy (mod reindexing) appearing in the support.
@@
Proposition 4.26.
Every $\beta$-support $S$ is an initial segment of some strong support.
Proof.
First, for each $(x,A)$ in the range of $S$ where $\delta={\tt min}(A_{1})$, $\gamma<{\tt min}(A_{2})$ and $x^{\circ}=f_{\gamma,\delta}(y,T)$, add the range of $T^{\uparrow A_{2}}$ to the end of $S$. Repeating this process $\omega$ times, we obtain a new support $S^{\prime}\supseteq S$ that satisfies the second property of being a strong support; the domain of $S^{\prime}$ is still a small ordinal as $\kappa$ is uncountable and regular. Finally, we append to $S$ all pairs ($(\{z\},A)$ where $z$ is in the interference of near-litters $M,N$ where $(M,A)$, $(N,A)$ are in the range of $S$. This produces a support $S^{\prime\prime}\supseteq S^{\prime}$ which is clearly strong. ∎
Definition 4.27 (coding function).
For any $(x,S)\in\tau_{\beta}^{+}$, we can define a function $\chi_{x,S}$ which sends $T=\pi[S]$ to $\pi(x)$ for every $T$ in the orbit of $S$ under the action of allowable permutations. We call such functions coding functions. Note that if $\pi[S]=\pi^{\prime}[S]$ then $(\pi^{-1}\circ\pi^{\prime})[S]=S$, so $(\pi^{-1}\circ\pi^{\prime})(x)=x$, so $\pi(x)=\pi^{\prime}(x)$, ensuring that the map $\chi_{x,S}$ for which we gave an implicit definition is well defined.
We will now define a certain combinatorial object that will characterize the orbit of a strong support under the action of allowable permutations (although we can define it even when the support in question is not strong).
Definition 4.28 (specification).
The specification of a $\beta$-support $S$ is the function $S^{*}$ with the same domain as $S$ given by the following criteria. We use the notation $S^{*}_{\epsilon}$ for $S^{*}(\epsilon)$.
  1. 1.
    If $S_{\epsilon}$ is ($(\{x\},A)$, then $S^{*}_{\epsilon}$ is given by
    $S^{*}_{\epsilon}=(0,A,\{\zeta:x\in\pi_{1}(S_{\zeta})\wedge A=\pi_{2}(S_{\zeta})\}).$
  2. 2.
    If $S_{\epsilon}$ is $(N,A)$ where $N$ is $A$-flexible, then $S^{*}_{\epsilon}$ is given by
    $S^{*}_{\epsilon}=(1,A,\{\zeta:(\exists N^{\prime}:S_{\zeta}=(N^{\prime},A)\wedge N\sim N^{\prime})\}).$
  3. 3.
    If $S_{\epsilon}$ is $(N,A)$ where $N$ is not $A$-flexible, so $|A|>2$, $\gamma={\tt min}(A_{1})$, and $N\sim f_{\delta,\gamma}(x,T)$ with $-1\leq\delta<{\tt min}(A_{2})$, then $S^{*}_{\epsilon}$ is given by
    $S^{*}_{\epsilon}=(2,A,\chi_{x,T},\{(\zeta,\eta):S_{\zeta}=(T^{\uparrow A_{2}})_{\eta}\}).$
@@
It should be evident that for any $\beta$-allowable permutation $\pi$ and $\beta$-support $S$, $(\pi[S])^{*}=S^{*}$. What is less evident and our first target result here is that if $S$ is a strong support then any strong $T$ with $T^{*}=S^{*}$ is the image of $S$ under the action of an allowable permutation: the specifications precisely code the orbits in the strong supports under the allowable permutations.
Lemma 4.30.
The specification of a strong $\beta$-support exactly determines the orbit in the action of $\beta$-allowable permutations on supports to which it belongs: if two strong $\beta$-supports have the same specification, they are in the same orbit.
Proof.
It is straightforward to see that if $S$ is a $\beta$-support and if $\pi$ is a $\beta$-allowable permutation, then $S^{*}=(\pi[S])^{*}$. The relationships between items in the support recorded in the specification are invariant under application of allowable permutations.
It remains to show that if $S$ and $T$ are strong supports, and $S^{*}=T^{*}$ is a specification for both, there is an allowable permutation $\pi$ such that $\pi[S]=T$.
We construct $\pi$ using the Freedom of Action Theorem (4.23). Define the $\beta$-action $\xi$ in the following way.
If $S_{\epsilon}=(M,A)$ for $M$ a near-litter, then $T_{\epsilon}=(N,A)$ for $N$ a near-litter, and we set $\xi_{A}(M)=N$. Note that if $S_{\delta}=(M,A)$ for some $\delta$, the fact that $S$ and $T$ have the same specification implies that $T_{\delta}=(N,A)$, ensuring that this is a well-defined function. Similarly, if $S_{\delta}=(M^{\prime},A)$ and $M^{\prime}\sim M$, then $T_{\delta}=(N^{\prime},A)$ for some $N^{\prime}\sim N$.
If we have $S_{\epsilon}=(\{x\},A)$, we will have $T_{\epsilon}=(\{y\},A)$ for some $y$, and so we will similarly set $\xi_{A}(x)=y$. Note that if $S_{\delta}=(\{x\},A)$, then $T_{\delta}=(\{y\},A)$.
As $S$ and $T$ are strong, this description indeed defines a $\beta$-action $\xi$. We now check that $\xi$ is coherent. If we assigned $\xi_{A}(M)=N$, then $(M,A)=S_{\epsilon}$ and $(N,A)=T_{\epsilon}$ for some $\epsilon$. As $S$ and $T$ have the same specification, $M$ is $A$-flexible if and only if $N$ is $A$-flexible.
Now suppose that $\xi_{A}(M)=N$ where $M^{\circ}=f_{\delta,\gamma}(x,U)$ and $\gamma={\tt min}(A_{1})$ and $\delta<{\tt min}(A_{2})$. Then $(M,A)=S_{\epsilon}$ and $(N,A)=T_{\epsilon}$ for some $\epsilon$. As $S$ and $T$ have the same specification, there is some $(y,V)\in\tau_{\delta}^{+}$ such that $N^{\circ}=f_{\delta,\gamma}(y,V)$ and
$\chi_{x,U}=\chi_{y,V};\quad\{(\zeta,\eta):S_{\zeta}=(U^{\uparrow A_{2}})_{\eta}\}=\{(\zeta,\eta):T_{\zeta}=(V^{\uparrow A_{2}})_{\eta}\}.$
Since $\chi_{x,U}=\chi_{y,V}$, there is a $\delta$-allowable permutation $\pi$ such that $\pi(x)=y$ and $\pi[U]=V$. So for any $(t,B)$ in the range of $U$, we have $S_{\zeta}=(t,B)$ if and only if $T_{\zeta}=(\pi_{B}(t),B)$. Therefore, as $S$ and $T$ are strong, we may conclude that $(\xi_{A_{2}})_{\delta}[U]=\pi[U]$, and that
$\xi_{A}(M)=N\sim f_{\delta,\gamma}(y,V)=f_{\delta,\gamma}(\pi(x),\pi[U])$
as required for coherence.
Finally, by the Freedom of Action Theorem (4.23), $\xi$ approximates some $\beta$-allowable permutation $\pi$, and in particular, $\pi$ satisfies $\pi[S]=T$, as required. ∎